# Membership Inference Attack against Large Language Model-based Recommendation Systems: A New Distillation-based Paradigm

### v1

@@ Introduction
Table 4: Comparative performance, training time and evaluation time of different types of attack models on all datasets. ”t-time” and ”e-time” represent attack model’s training time and evaluation time respectively. The best results on each dataset are in bold.
Movie Ours Vector Entropy Confidence Loss
acc% 80.00 59.38 58.75 69.37 78.75
@@
Deli Ours Vector Entropy Confidence Loss
acc% 64.75 55.48 61.97 60.31 63.84
@@
Table 5: Comparison of the attack performance between our fused features and individual features on MovieLens and Delicious. ”Movie” refers to MovieLens and ”Deli” refers to Delicious.
Last.FM ours /vector /entropy /confidence /loss
acc% 79.55 70.45 72.73 59.09 54.55
@@
Book ours /vector /entropy /confidence /loss
acc% 86.60 85.31 84.79 84.28 85.57
@@
Table 6: Comparison of the performance between our fused feature and that after removing each feature respectively on Last.FM and Book-Crossing. ”/*” means removing ”*” from fused feature. ”Book” refers to Book-Crossing.
As shown in Figure 4, we found that the impact of different $\alpha$ values on the attack performance doesn’t show a completely consistent trend across datasets. For instance, on Last.FM, the attack effectiveness fluctuates significantly with varying $\alpha$ values, and on MovieLens, the optimal attack performance is achieved when $\alpha$ increases to 0.9. For Book-Crossing and Delicious, the best results occur when $\alpha$ equals 1. Therefore, in practical applications, selecting an appropriate $\alpha$ value according to the specific dataset can achieve optimal attack performance. Although the optimal values of $\alpha$ vary across different datasets, by changing it, our paradigm can achieve superior attack performance compared to random attack in the four datasets.
Attack Model Type Selection. We analyzed the impact of the complexity of the attack model on attack perfromance. We selected a simple logistic classifier and a relatively complex model architecture for training the attack model. To be precise, we choose Bert (Devlin et al. 2019) to investigate the impact of attack model complexity on both attack effectiveness and time cost.
Effect of Fused Features. To evaluate the effectiveness of fused features, we compared the attack performance of our attack model with those trained using individual features (Table 5) and those trained after removing each feature (Table 6). In Table 5, our method outperforms attack models trained with individual features in all situations, indicating that fused feature can accumulate the differences in individual features between member and non-member data. In Table 6, our method also surpasses attack models trained after removing any individual feature, indicating that each feature in the fused feature contributes to improving the attack performance.
The results also demonstrate that different features contribute variably to the attack effectiveness. Among the four features we employed, loss exhibits the highest contribution to the attack performance, whereas vector shows the minimal enhancement effect. Due to space limitations, results in Table 5 and Table 6 on the rest datasets are shown in the additional resources.
Conclusion and Future work
In this paper, we propose a novel MIA paradigm against LLM-based recommendation systems. We obtain reference model using knowledge distillation to enhance the distinction between member and non-member data. We expand natural language datasets in recommendation field and generate stronger fused feature for attack model. Comprehensive experiments on various datasets demonstrate the superiority of our paradigm.
Currently our paradigm operates under gray-box background knowledge for MIA. Black-box distillation may provide a more rigorous attack premise. Our extended dataset only contains user interaction histories without user preferences, resulting in less semantically rich data. In future research, we will explore effective methods to extract higher-quality semantic information to enhance attack performance on semantically sparse datasets. We will also adapt our MIA paradigm to other LLM-based recommendation tasks to prove the scalability of our paradigm.
Appendix
Additional Experiments
Effect of Fused Features. We supplement additional results that were omitted from Tables 5 and 6 in the main text due to space limitations. The comparative attack performance of our attack model and those trained using individual features on Last.FM and Book-Crossing are shown in Table 7. The comparative attack performance of our attack model and those trained after removing each feature on MovieLens and Delicious are shown in Table 8.
Last.FM
ours vector entropy confidence loss
acc 79.55 50.00 45.45 65.91 52.27
recall 95.45 45.45 40.91 72.73 95.00
f1 82.35 47.62 42.86 68.09 66.67
auc 77.48 46.90 50.00 70.87 72.73
Book-Crossing
ours vector entropy confidence loss
acc 86.60 51.29 84.54 81.19 85.05
recall 98.99 52.58 95.88 98.45 98.97
f1 88.02 51.91 86.11 83.96 86.88
auc 96.71 53.71 93.54 94.14 95.34
Table 7: Comparison of the attack performance between our fused features and that with individual features on Last.FM and Book-Crossing.
MovieLens
ours /vector /entropy /confidence /loss
acc 80.00 80.63 79.37 80.00 69.37
recall 96.25 92.50 90.00 91.25 75.00
f1 82.80 82.68 81.36 82.02 71.01
auc 90.84 86.36 85.95 85.27 75.42
Delicious
ours /vector /entropy /confidence /loss
acc 64.75 64.23 63.01 63.49 62.27
recall 88.81 86.80 85.87 86.12 80.32
f1 72.28 71.52 70.61 70.94 68.78
auc 67.55 68.29 66.03 66.83 64.48
@@
Table 8: Comparison of the performance between our fused feature and that after removing each feature respectively on MovieLens and Delicious. ”/*” means removing ”*” from fused feature.
Con Wei_Con Sum Wei_Sum MLP Wei_MLP Com Wei_Com
acc 80.63 86.00 76.88 80.00 83.75 85.00 76.25 78.75
recall 92.50 95.00 90.00 97.50 91.25 93.75 91.02 93.20
f1 82.68 87.86 79.56 82.98 84.88 86.21 80.22 81.52
auc 89.89 90.87 80.95 87.75 91.38 90.14 88.06 87.73
@@
Table 9: Comparison of the performance of different feature fusion strategies on MovieLens. ”Con” refers to ”Concatenation”, ”Wei_*” refers to ”Weighted *” and ”com” refers to ”MLP Compression”, The best results are in bold.
Figure 5: Data sample demonstration.
@@
Feature Fusion Strategy. We analyzed the impact of different feature fusion strategies on attack performance. Our analysis considers the influence of feature weight and dimension on the attack performance. In particular, we consider the dimensional differences in the features we use. The fusion methods that that we adopted include:
@@
  • Sum: Add confidence, entropy and loss as a feature and then concatenate it with vector feature.
@@
  • Weighted MLP: Based on MLP method above, we perform weighted concatenation of the four vectors to obtain a 4*512D weighted vector.
  • MLP Compression: Based on MLP method, we remap the obtained 4*512D vectors back to 512 dimensions to avoid unnecessary computational costs caused by excessively long vectors.
  • Weighted MLP Compression: Based on MLP Compression method, we perform weighted concatenation of the four vectors to obtain a 512D weighted vector. The weights are determined by the importance of each feature verified in the multi-feature fusion ablation experiments(Table 5(in main text), Table 6(in main text), Table 7 and Table 8). We set 0.5, 0.3 and 0.2 as original weight for loss, confidence and vector respectively.
@@
Data Illustration
As shown in Figure 5, we illustrate data samples of the datasets we adopted from existing work (Book-Crossing and MovieLens) and our extended datasets(Last.FM and Delicious).

### v2

@@ Introduction
As shown in Figure 4, we found that the impact of different $\alpha$ values on the attack performance does not show a completely consistent trend across datasets. For instance, on Last.FM, the attack effectiveness fluctuates significantly with varying $\alpha$ values, and on MovieLens, the optimal attack performance is achieved when $\alpha$ increases to 0.9. For Book-Crossing and Delicious, the best results occur when $\alpha$ equals 1. Therefore, in practical applications, selecting an appropriate $\alpha$ value according to the specific dataset can achieve optimal attack performance. In fact, the determination of $\alpha$ is influenced by various factors, including the model architecture and dataset characteristics, as the implementation focus varies depending on the specific data and task. Although the optimal values of $\alpha$ vary across different datasets, by changing it, our paradigm can achieve superior attack performance compared to a random attack in the four datasets.
Figure 4: The impact of different $\alpha$ values on the performance of our paradigm on the four datasets respectively.
@@
Table 6: Comparative performance, training time and evaluation time of different types of attack models on all datasets. “t-time” and “e-time” represent the attack model’s training time and evaluation time respectively. The best results on each dataset are in bold.
Data Metrics Ours Vec Entr Conf Loss
LF acc% 79.55 50.00 45.45 65.91 52.27
recall% 95.45 45.45 40.91 72.73 95.00
f1% 82.35 47.62 42.86 68.09 66.67
auc% 77.48 46.90 50.00 70.87 72.73
ML acc% 80.00 59.38 58.75 69.37 78.75
@@
BC acc% 86.60 51.29 84.54 81.19 85.05
recall% 98.99 52.58 95.88 98.45 98.97
f1% 88.02 51.91 86.11 83.96 86.88
auc% 96.71 53.71 93.54 94.14 95.34
DL acc% 64.75 55.48 61.97 60.31 63.84
@@
Table 7: Comparison of the attack performance between our fused features and individual features on the four datasets. “Vec”, “Entr” and “Conf” mean Vector, Entropy and Confidence respectively.
Attack Model Type Selection. We analyzed the impact of the complexity of the attack model on attack perfromance. We selected a simple logistic classifier and a relatively complex model architecture for training the attack model. To be precise, we choose BERT (devlin2019bert) to investigate the impact of attack model complexity on both attack effectiveness and time cost.
Effect of Fused Features. To evaluate the effectiveness of fused features, we compared the attack performance of our attack model with those trained using individual features (Table 7) and those trained after removing each feature (Table 8). In Table 7, our method outperforms attack models trained with individual features in most situations, indicating that fused feature can accumulate the differences in individual features between member and non-member data. In Table 8, our method also surpasses most attack models trained after removing any individual feature, indicating that each feature in the fused feature contributes to improving the attack performance.
The results also demonstrate that different features contribute variably to the attack effectiveness. Among the four features we employed, loss exhibits the highest contribution to the attack performance, whereas vector shows the minimal enhancement effect.
Data Metrics Ours /Vec /Entr /Conf /Loss
LF acc% 79.55 70.45 72.73 59.09 54.55
@@
BC acc% 86.60 85.31 84.79 84.28 85.57
@@
ML acc% 80.00 80.63 79.37 80.00 69.37
recall% 96.25 92.50 90.00 91.25 75.00
f1% 82.80 82.68 81.36 82.02 71.01
auc% 90.84 86.36 85.95 85.27 75.42
DL acc% 64.75 64.23 63.01 63.49 62.27
recall% 88.81 86.80 85.87 86.12 80.32
f1% 72.28 71.52 70.61 70.94 68.78
auc% 67.55 68.29 66.03 66.83 64.48
@@
Table 8: Comparison of the performance between our fused feature and that after removing each feature respectively on the four datasets. “/*” means removing “*” from the fused feature.
Con Wei_Con Sum Wei_Sum MLP Wei_MLP MLP_Com Wei_MLP_Com
acc% 80.63 86.00 76.88 80.00 83.75 85.00 76.25 78.75
recall% 92.50 95.00 90.00 97.50 91.25 93.75 91.02 93.20
f1% 82.68 87.86 79.56 82.98 84.88 86.21 80.22 81.52
auc% 89.89 90.87 80.95 87.75 91.38 90.14 88.06 87.73
@@
Table 9: Comparison of the performance of different feature fusion strategies on MovieLens. “Con” means “Concatenation”, “Wei” means “Weighted” and “Com” means “Compression”. The best results in weight inclusion comparisons are in bold.
Figure 5: Data sample demonstration of the four datasets.
@@
Feature Fusion Strategy. We analyzed the impact of different feature fusion strategies on attack performance. Our analysis considers the influence of feature weight and dimension on the attack performance. In particular, we consider the dimensional differences in the features we use. The fusion methods that we adopted include:
@@
  • Sum: Add confidence, entropy and loss as a feature and then concatenate it with the vector feature.
@@
  • Weighted MLP: Based on the MLP method above, we perform weighted concatenation of the four vectors to obtain a 4*512D weighted vector.
  • MLP Compression: Based on the MLP method, we remap the obtained 4*512D vectors back to 512 dimensions to avoid unnecessary computational costs caused by excessively long vectors.
  • Weighted MLP Compression: Based on the MLP Compression method, we perform weighted concatenation of the four vectors to obtain a 512D weighted vector. The weights are determined by the importance of each feature verified in the multi-feature fusion ablation experiments.
@@
Data Illustration. Based on the existing MovieLens and Book-Crossing datasets, we constructed natural language dataset formats for Last.FM and Delicious using the same prompt structure. Items were selected in chronological order of user history, with the most recent items used to determine user preferences and the remaining items serving as user history. Since Last.FM and Delicious only contain user historical records without explicit positive/negative item preferences (unlike MovieLens and Book-Crossing), the constructed data from Last.FM and Delicious contains less effective information than that from MovieLens and Book-Crossing. This may explain why in some experimental results, the data volume of these two datasets does not reach the expected attack effectiveness. As shown in Figure 5, we illustrate data samples of the datasets we adopted from existing work (Book-Crossing and MovieLens) and our extended datasets (Last.FM and Delicious).
Conclusion and Future Work
In this paper, we propose a novel MIA paradigm against LLM-based recommendation systems. We obtain a reference model using knowledge distillation to enhance the distinction between member and non-member data. We expand natural language datasets in recommendation field and generate stronger fused feature for attack model. Comprehensive experiments on various datasets demonstrate the superiority of our paradigm.
Our paradigm operates under gray-box so far. Black-box distillation may provide a more rigorous attack premise. Our extended dataset only contains user interaction histories without user preferences, resulting in less semantically rich data. We will explore methods to extract higher-quality information to enhance attack performance on semantically sparse datasets. We will also adapt our method to other LLM-based recommendation tasks to prove its scalability.
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under Grant (2023YFC3310700), the National Natural Science Foundation of China (62572040, 62202041), the Beijing Natural Science Foundation (JQ24019), and the Fundamental Research Funds for the Central Universities (2025JBMC011).